第45回環論および表現論シンポジウム (2012年)

第二報

第 45 回環論および表現論シンポジウムのプログラムが出来ましたのでお知らせいたします。また懇親会もあわせてご案内申し上げます。このシンポジウムは以下の研究費の援助の元に行われます。

日本学術振興会科学研究費補助金基盤研究 (C)

研究代表者 : 花木章秀 (信州大学) 研究代表者 : 馬場良始 (大阪教育大学) 研究代表者 : 池畑秀一 (岡山大学) 研究代表者 : 河田成人 (大阪市立大学) 研究代表者 : 西中恒和 (岡山商科大学) 研究代表者 : 小池寿俊 (沖縄高専)

プログラム責任者小池寿俊 (沖縄高専)

記

日程 : 2012 年 9 月 7 日 (金)〜9 月 9 日 (日) 場所 : 信州大学理学部第 1 講義室

〒 390-8621 長野県松本市旭 3-1-1

※駅から会場までのバスの本数が限られておりますのでご注意下さい。 http://www.alpico.co.jp/access/matsumoto/ http://www.alpico.co.jp/access/matsumoto/pdf/shindaiyokota20120316.pdf

「信州大学前」または「大学西門」下車

会場責任者 : 花木章秀 (信州大学 hanaki@shinshu-u.ac.jp) プログラム責任者 :小池寿俊 (沖縄高専 koike@okinawa-ct.ac.jp) 懇親会 : 日時 : 9 月 8 日 (土) 18 時 40 分〜

会場 : 信州大学旭会館食堂 (シンポジウム会場から徒歩１分)

会費 : 4,000 円懇親会参加希望者は 8 月 27 日 (月) までに、 E-mail にて、会場責任者・花木章秀氏

(hanaki@shinshu-u.ac.jp) までお知らせ下さい。多数の方のご参加をお待ちしています。

9 月 7 日 (金)

9:00–9:30 Monika Rianti Helmi (Andalas University), 丸林英俊 (徳島文理大学),

植田玲 (島根大学) Skew Rees rings which are maximal orders

9:40–10:10 亀山統胤 (信州大学), 星野光男 (筑波大学), 古賀寛尚 (筑波大学)

Constructions of Auslander-Gorenstein local rings 10:20–10:50 上山健太 (静岡大学)

Noncommutative graded Gorenstein isolated singularities 11:00–11:30 松田一徳 (名古屋大学)

Characterization of Gorenstein strongly Koszul Hibi rings by F-invariants 11:40–12:10 東谷章弘 (大阪大学)

Toric rings arising from cyclic polytopes 13:30–14:00 神田遼 (名古屋大学)

Classifying Serre subcategories via atom spectrum 14:10–14:40 源泰幸 (名古屋大学)

Derived Gabriel topology, localization and completion of dg-algebras 14:50–15:20 竹花靖彦 (函館工業高等専門学校)

A generalization of Goldie torsion theory 15:40–16:10 本瀬香

Power residues 16:20–16:50 山中聡 (岡山大学)

On separable polynomials in skew polynomial rings 17:00–17:30 倉富要輔 (北九州工業高等専門学校)

On Goldie extending modules with finite internal exchange property

9 月 8 日 (土)

9:00–9:30 木村真弓 (静岡大学), 浅芝秀人 (静岡大学) Derived equivalence classification of generalized multifold extensions of piecewise hereditary algebras of tree type

9:40–10:10 Liu Yu (名古屋大学)

Abelian subquotients of exact categories by rigid subcategories 10:20–10:50 荒谷督司 (徳山工業高等専門学校)

Dimensions of triangulated categories with respect to subcategories 11:00–11:30 吉田健一 (日本大学), 渡辺敬一 (日本大学)

Ulrich modules and Special modules over 2-dimensional rational singularities 11:40–12:10 木村嘉之 (大阪市立大学)

Quiver varieties and quantum cluster algebras 13:30–14:00 小西正秀 (名古屋大学)

Cyclotomic KLR algebras of cyclic quivers 14:10–14:40 Luo Xueyu (名古屋大学)

Realizing cluster categories as stable categories of lattices (A_{n} case) 14:50–15:20 Gustavo Jasso (名古屋大学)

Cluster-tilted algebras of canonical type and quivers with potentials 15:40–16:10 伊山修 (名古屋大学)

τ-tilting theory (仮題) 16:20–16:50 足立崇英 (名古屋大学)

τ-tilting modules for self-injective Nakayama algebras 17:00–17:30 Martin Herschend (名古屋大学)

2-hereditary algebras and quivers with potential 17:40–18:10 水野有哉 (名古屋大学)

Selfinjective algebras and quivers with potentials 18:40– 懇親会

9 月 9 日 (日)

9:00–9:30 植松盛夫 (上武大学)

On the relation of the upper bound of global dimension and the length of serial algebra which has finite global dimension

9:40–10:10 Laurent Demonet (名古屋大学)

TBA 10:20–10:50 Erik Darpö (名古屋大学)

On the representation rings of the dihedral 2-groups 11:00–11:30 小原大樹 (東京理科大学), 古谷貴彦 (東京理科大学)

On Hochschild cohomology of a class of weakly symmetric algebras with radical cube zero

11:40–12:10 古谷貴彦 (東京理科大学)

On the existence of a bypass in the Auslaner-Reiten quiver